The concentration of an aqueous solution of sodium hydroxide is prepared by weighing 5.0 grams of solid NaOH (MW = 40.0 grams per mole) using an analytical balance with a standard deviation of 0.001 gram. The solid is then dissolved in enough distilled water to the full volume of a 1-L volumetric flask whose standard deviation is 5.0 milliliters. What is the molar concentration of the NaOH solution?.


1
Expert's answer
2021-02-25T03:40:48-0500

Mass given = 5.0g

Molecular Weight= 40g/mol

Volume= 1L


Assuming that the deviations are insignificant


No of moles of NaOH = 5/40 = 0.125mol


Molar concentration = Amount of substance/ Volume

= 0.125/1= 0.125mol/L


Hence, the Molar concentration of the NaOH solution is 0.125mol/L